OpenStudy (lokiando):

Easy question! Which is a difference of cubes? a) x^6 – 27 b) x^15 – 36 c) x^16 – 64 d) x^5 – 125

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (lokiando):

this is an example

OpenStudy (lokiando):

c

OpenStudy (lokiando):

x^16-64

OpenStudy (skullpatrol):

why?

OpenStudy (lokiando):

8*8=64

OpenStudy (skullpatrol):

how does that fit the pattern of being a "difference of cubes"?

OpenStudy (skullpatrol):



OpenStudy (lokiando):

its D

OpenStudy (skullpatrol):

why?

OpenStudy (lokiando):

because 5^3=125

OpenStudy (skullpatrol):

prove it by writing it as a difference of cubes

OpenStudy (skullpatrol):

you are right 5^3 = 125, but what about x^5?

OpenStudy (lokiando):

0

OpenStudy (skullpatrol):